There are 6!/3! = 120 possible combinations of marble colour. Of these, only 2 are "good". This gives a probability of 2/120 = 1/60 = 0.0167, or about one and two-thirds of a percent.
Under what circumstances might conscription be necessary?
I assume your question refers to The United States in our current circumstances, i.e. the War on Terror. I come from the Cold War generation that faced a universal draft. In fact, I volunteered for the army because I knew if I didn't I'd be drafted. I served through the Cuban Missile Crisis and the Berlin Wall. I was serving when President Kennedy was shot and we weren't sure whether there was a conspiracy or if the Soviets were involved. We didn't stop the draft until 1973, just two years before we finally gave up and got out of Vietnam.
Under what circumstances might conscription be necessary? Whether you agree with the war in Iraq or not, the bottom line is that our all-volunteer military is stretched perilously thin in Iraq and Afghanistan. We have not been able to capture Osama bin Laden or any of his higher ups. The War in Iraq has nearly sucked dry the energy of all the Armed Forces. We are extending tours for our soldiers, and some of our troops have done three tours so far in either Iraq or Afghanistan, with no end in sight. The troops we have are becoming exhausted, and enlistment rates are down, especially for the army.
Despite the Pentagon's own admission that our self sacrificing volunteer troops are nearing the breaking point, the present administration is rattling the saber at both Iran and North Korea over nuclear weapons and other issues, especially the mounting evidence that Iran is providing deadly weapons and training to terrorists in Iraq. Should we somehow end up in a war with either or both, where would the soldiers come from? It's at that point that I think conscription might be necessary. It could even become necessary should enlistment rates fall to the point that we can no longer supply the troops to keep up the "surge" in Iraq, if the next administration chooses to continue the fight. If the all volunteer force is too small and depleted to do the job, history shows that the government is likely to turn to conscription. It did in the Civil War and both World Wars, not to mention The Cold War and Vietnam.
The really big question is, what happens if the government tries to reinstate the draft? How many young people would obey the law? How many would burn their draft cards or flee? Do we conscript women? Should women fight? (They do now.) Do we have to agree with a war in order to fight it? Can those who don't agree say No? These are all questions I personally hope we never have to answer.
"This relationship isn't working for me. I need more actions, that shows your interested. I feel like your not really interested, so I'm ending it."
You need to say it to his face. Edit the wording where needed. But be confident. This is what you want and what you need and what your not getting. Don't point your finger at him and tell him what all he did wrong. Tell him what you needed and didn't get. Resolve it short and sweet.
You can ammend this statement with, "I like you and I'd like to be in this relationship with you. If that's what you want too, then you have a few hours/days to think about it before I have to end it."
Don't be afraid to have a friend standing near by to get you out of a bad situation. If his anger gets explosive you want someone that can help you to get out of there. But keep them out of ear shot, this is a private matter between you and him, your friend should only be close enough to see if he puts his hands on you.
Give example of variable that is normally distributed?
Some examples of anormal distribution are: heights of men and women, weights of men and women, IQ's, body temperature, and many manufacturing processes to name a few.
If you roll a regular 6-face die what is the probibility of rolling a 2?
There are 6 outcomes when a die is rolled; 2 is one outcome so the probability of rolling a 2 is 1/6.
What are the possible outcomes of a coin that is flipped?
The possible outcomes of a coin that is flipped are heads or tails.
The mode and standard deviation of a distribution are 55 and 4.33 respectively?
The statement is probably: The mean and standard deviation of a distribution are 55 and 4.33 respectively.
It is 0.8181... recurring.
15
It should make marketing easier because the amount purchased has a smaller variance. This means that the numbers are closer to 12 and so there is less of a need to carry large stocks for when the demand is high.
"[Unoka] was poor and his wife and children had barely enough to eat."
"But [Okonkwo's] whole life was dominated by fear, the fear of failure and of weakness. . . . It was the fear of himself, lest he should be found to resemble his father. . . ."
"Unoka loved . . . the first kites that returned with the dry season, and the children who sang songs of welcome to them."
"If any money came [Unoka's] way . . . he immediately bought gourds of palm-wine, called round his neighbors and made merry."
"Okonkwo is a strong, disciplined man but he cannot enjoy simple things in life." "And so Okonkwo was ruled by one passion - to hate everything that his father Unoka had loved."
Is the probability 1.2 nil possible?
Probabilities go from 0 to 1. This is a steady increase with 0.5 meaning 50% of "equally likely to happen as not." 0 probability means that it is impossible something will happen. 1 probability means it is definite something will happen. Therefore, you cannot have probability 1.2 as this is more than "definitely going to happen."
(85 - 58)/11 = 27/11 = 2.4545.. sd
What is the probability of choosing a card at from a pack of cards and NOT picking a red jack?
It is 50/52 or 0.9615
pons connect the brain and the central nervous system with a transverse band of nerves. they are helpful in controlling the respiration and moving the head as per audiovisual movements.
How might a person use probability in his or her job?
There are a lot of ways a person can use probability in his/her job. For example, People who bet on horse races use the odds and their knowledge to determine which horse will have the best chance of win,place,show Another example are poker players, quite often, you will see a percentage in the corner of your screen if you watch poker on TV
very nice imature ryme an so you 1 in 3 guys you ment o school with and kno now are gay.
The probability of tossing a coin and getting a tail and rolling a dice and getting a four?
It is (1/2)*(1/6) = 1/12
When doing a t-test what does normal distributed in population mean?
It means that the random variable of interest is Normally distributed and so the t-distribution is an appropriate distribution for the test rather than just an approximation.
What is the probability of rolling a square number on a fair sided dice?
Assuming that it is a 6-sided die, the answer is 1/3.